# Evolutionary Retrosynthetic Route Planning

### v1

@@ III Proposed method
This section initiates by framing the retrosynthetic problem as a tree search problem, where the objective is to find a sequence of routes leading from the root node to leaf nodes. Then the evolutionary algorithm efficiently solves the tree search problem.
@@ III-A Retrosynthetic problem
where $R$ and $W$ represent reactants, $P$ represents product in Eq 1a.
@@
Figure 4: K-cross search tree structure transformed by multi-step retrosynthesis process. Gray root node represents target product, $NN$ represents transformer model, blue leaf node stands for raw material in the building block dataset.
@@ III-B Retrosynthetic optimization model
The number of potential chemical structures for organic molecules can reach an enormous scale, up to about $10^{60}$ compounds[58]. However, only a much smaller subset of compounds exhibits reasonable structural characteristics. Therefore, constructing a search space that explores the natural distribution of organic structures with structural rationality is a crucial challenge. The search space is defined as the top-k reactants with the highest inferred confidence in the single-step retrosynthetic reaction. The choice of the value $k$ is crucial as it impacts the search efficiency and the quality of the solutions. If the search space is too small, there is a risk of missing the global optimal solution. Conversely, if the search space is too large, the algorithm will consume more GPU memory and time, potentially leading to local optimality and reduced efficiency. Due to the vast search space, tree search alone may not be sufficient to effectively solve the problem. Therefore, heuristic algorithms are employed to address this limitation and provide better solutions.
To tackle the retrosynthetic problem, it is converted into an optimization problem resembling tree search. Suppose that a multi-step retrosynthesis involves $q$ single-step retrosynthesis processes. We use $(x,r,\beta)$ to represent a multi-step retrosynthesis, where $x=(x^{1},x^{2},\cdots,x^{q})$, and $x^{i}\in\{1,2,\cdots,k\}$ represents the $x^{i}$-th possible output of the $i$-th single-step retrosynthesis. Let $r=(r^{1},r^{2},\cdots,r^{q})$ denotes decoding results from $x$ using single-step retrosynthesis, as shown in Fig. 3, and $r^{i}$ represents SMILES expression (eg.,Fc1cccc(F)c1CBr.[N-]=[N+]=[N-]) of the $i$-th single-step retrosynthesis. $\beta=(\beta^{1},\beta^{2},\cdots,\beta^{q})$ denotes the corresponding probability of $r$. Take $(<2,2,1>,<r^{1},r^{2},r^{3}>,<0.8,0.9,0.7>)$ as an example, there are 3 steps: in the 1st step, $r^{1}$ is chosen as the second reactant of $NN$ with a probability of $\beta^{1}$=0.8; in the 2nd step, $r^{2}$ is chosen as the second reactant of $NN$ with a probability of $\beta^{2}$=0.9; in the 3rd step, $r^{3}$ is chosen as the first reactant of $NN$ with a probability of $\beta^{3}$=0.7 and it is a known molecule in $\Psi$.
@@ III-C2 Variation
The retrosynthetic problem is tackled using the Estimation of Distribution Algorithm (EDA) [61] during the execution of the EA process. EDA operator stands out among other operators as it generates new solutions by sampling from a histogram probabilistic model. This characteristic enables EDA to produce better or more similar solutions and converge faster. The basic framework of the proposed method is depicted in Fig. 6. The process of variation consists the establishment of the probability model and sample. To construct the probability model, the variation space $[a^{i,0},a^{i,M}]$ for the $i$-th variable is divided into M bins, where $a^{i,0}=0$ and $a^{i,M}=1$ are the boundaries of the bins. Subsequently, $a^{i,1}$ and $a^{i,M-1}$ can be set as follows:
@@ IV Experiment
In this section, we will begin by introducing the datasets and parameter settings. Next, we will conduct a comparative study using various methods and perform extensive experiments on four case products. Finally, we have provided several charts and conducted corresponding analysis.
@@ IV-A Datasets
In this experiment, we performed retrosynthetic route planning for four case products. Two commonly used benchmark datasets, namely USPTO_50K [62, 63] and USPTO_MIT, were utilized to train single-step retrosynthetic model. The USPTO_50K dataset was extracted from the patent literature of the United States, while USPTO_MIT was previously employed by Lin et al.[3]. The distribution of reaction classes in the two datasets is presented in Table I. To define the terminal nodes or reactants, the building block dataset [3] is built, which contains 93563 commercially available molecules.
TABLE I: Descriptions of ten reaction classes and the ratio of USPTO_50K and USPTO_MIT
@@
IV-B Parameter settings
The default population size and thread pool size were both set to 42. The search process terminated after 200 iterations. For this optimization, the number of bins $M$ was set to 10. The boundaries of each dimension in the individual ranged from 0 to 1. The experiments in this paper were conducted with 30 independent runs, and the mean and variance were used for result analysis. The running program is written using python scripts. During the search process, the retrosynthetic route was explored using 3 NVIDIA GeForce RTX 4090 GPUs.
TABLE II: Analysis on the number (mean(std.dev.))[percentage] of calling single-step model for MCTS and EA. Percentage in “[$\cdot$]” expresses increase or decrease of EA compared with MCTS in the same column and under the same product. over 30 independent runs.
@@ IV-C Comparative experiment
In the experimental section, we conducted a series of comparative experiments to validate the effectiveness of the proposed algorithm. In particular, we scrutinized the impact on the number of calling single-step model, the search capability in a larger search space, and the count of feasible search solutions. Additionally, we designed several objective optimization functions and tested the algorithm’s search time.
@@ IV-C1 Comparison experiment on calling single-step model
The frequency of calling single-step model directly affects both the search time and search efficiency. MCTS and EA were employed to search for one, two, and three solutions for the four case products, as depicted in Table II. The values in the table indicate the average and standard deviation of the number of calling single-step model. As the number of solutions increase, both MCTS and EA require a higher count of model calls. However, EA consistently outperforms MCTS for one, two, and three solutions. Furthermore, across the four case products, the number of calling single-step model reduced by an average of 78.6%, 59.0%, and 53.9% respectively. Consequently, EA exhibits faster search speed and higher search efficiency when compared to MCTS.
@@ IV-C2 Comparison experiment on the search capability
To compare the search capabilities of the two algorithms in different search space, we conducted tests with three different beam sizes: top-10, top-15, and top-20.
Firstly, with the expansion of search space, the number of calling single-step model for both MCTS and EA increases significantly. As shown in Table IV(a), due to the advantage of EA search mechanism, under the beam size of top-10, top-15, and top-20, the number of calling single-step model of EA decreased by 53.9%, 27.6%, and 16.8% on average, respectively, compared to that of MCTS based on searching 3 feasible solutions.
Secondly, this experiment is conducted with a beam size of top-10. Notably, there is a progression in the complexity of molecular structures from DemoA to DemoD, which subsequently leads to longer search time. When focusing on the same target product, the search time for EA decreased significantly compared to that of MCTS, with an average reduction of 83.9% across the four case products based on searching 3 feasible solutions, as shown in Table IV(b).
Thirdly, as the search space increased in Table IV(c), the complexity of the search also rises, meanwhile, the number of search routes for four case products gradually decreased under the same number of iterations. However, EA outperforms MCTS in terms of the number of search routes across top-10, top-15, and top-20. Among them, the performance on top-15 and top-20 is normal, but the performance on top-10 is particularly noteworthy, with the number of search routes increasing 5.13 times on average. This could be attributed to the relatively simplistic and easily searchable molecular structure of DemoA.
@@
TABLE III: Analysis on search capability (mean(std.dev.))[percentage] on beam size top-10, top-15, top-20. Percentage in “[$\cdot$]” expresses increase or decrease of EA compared with MCTS in the same column and under the same product. Data in the table is over 30 independent runs.
@@
(a) Num of calling single-step model based on searching 3 feasible solutions.
@@
(b) Search time (sec) based on searching 3 feasible solutions.
@@
DemoA EA 8.47e+1(2.08e+1)[2017$\uparrow$] 3.86e+0(1.01e+0)[18.0$\uparrow$] 2.71e+0(1.67e+0)[1.9$\uparrow$]
@@ IV-C3 Comparison experiment on different objective functions
$f_{1}$ $=-f(x_{i})$ (10)
$f_{2}$ $=-e^{f(x_{i})}$ (11)
$f_{3}$ $=-\frac{f(x_{i})}{\sum_{j=1}^{n}f(x_{j})}$ (12)
During the selection process, solutions are ordered based on the magnitudes of their fitness values for $f_{1}$ and $f_{2}$, with $f_{1}$ is the negation of the original objective function, and $f_{2}$ engineered to increase the separation between target values. Meanwhile, $f_{3}$ employs the roulette wheel selection[64] sampling method, allowing for the selection of offspring solutions according to the likelihood that an individual can be sampled. Fig. 8 shows that $f_{1}$ yields more feasible solutions compared to the other objective functions, making it the most effective among them. To further assess the convergence of the three different objective functions on the four case products, we normalize the value range of $f_{2}$ to match the same range as $f_{1}$ and $f_{3}$, which is from -1 to 0. As shown in Fig. 9, the results clearly demonstrate that $f_{1}$ outperforms the other two objective functions on DemoA, DemoB, and DemoC. Although $f_{3}$ shows some similarity to $f_{1}$ on DemoD, $f_{1}$ still performs better after 80-th iteration.
@@
Figure 9: Convergence curves of three different objective functions. Objective function $f_{1}$ outperforms the other two functions on the four target products.
@@ IV-D Visualization of results
In the retrosynthetic analysis, we used four case products that were not present in the training dataset. DemoA refers to Rufinamide[3], an antiepileptic triazole derivative. DemoB, DemoC, and DemoD are from Klucznik et al.[27], Li et al.[65], and Segler et al.[1] respectively. The retrosynthesis of these four products required 4, 2, 3, and 6 steps, respectively. As depicted in Fig. 11, DemoA’s search procedure comprises four steps leading to the final products being in $\Psi$. The impacted functional groups are highlighted in red. The digits before and after the “.” indicate the reaction type and the substance’s positioning in the results of the single-step model inference, correspondingly. These four case products have received approval from chemical experts.
During the search process, we visualized the distribution of the encoded population, as shown in Fig. 12. As the iteration increases, the population distributions gradually converge to the similar position. DemoA and DemoB demonstrate nearly rapid convergence after approximately 20 iterations, which can be attributed to their relatively simple molecular structures. In contrast, DemoC and DemoD, with more intricate molecular configurations, necessitate a broader range of reaction rules for analysis, presenting greater search challenges. As a result, these compounds converge around the 60-th iteration. Despite the intricate nature of the problem and the gradual rate of convergence, all four compounds ultimately achieve convergence.
@@
EA consistently outperforms MCTS in different aspects, as shown in Fig. 13. In the context of identifying three feasible solutions, there is a considerable decrease in the number of calling single-step model and search time. Additionally, particularly noteworthy is EA’s significant enhancement in searching feasible solutions under 200 iterations, especially for DemoA, because of its relatively simple molecular structure. Furthermore, the parallel implementation of EA results in an exponential reduction in time, making it a highly efficient approach for retrosynthetic route planning. Overall, the utilization of EA in retrosynthetic route planning demonstrates its capability to efficiently handle complex and challenging synthesis problems, thereby providing chemists with valuable and promising potential routes for target compounds.
Figure 13: The experiment of calling single-step model and search time is conducted based on identifying 3 feasible solutions. Additionally, the experiment of the number of feasible routes is carried out based on 200 iterations.
@@ V Conclusion
This study presents a novel and practical method for retrosynthesis of compounds, utilizing Evolutionary Algorithms (EA) for the first time to address the retrosynthetic problem. By encoding the complex retrosynthesis into an optimization problem, the retrosynthetic process is effectively simplified. This approach allows for a well-defined search space and the establishment of a suitable objective function. At the same time, the use of parallel computation further boosts the search efficiency. According to the experimental results, the EA demonstrates a remarkable reduction in the number of calling single-step model, averaging 53.9%, when compared to the MCTS based on 3 feasible solutions. Additionally, the search time is reduced by an average of 83.9%, and the number of search routes increases by 5 times. The results demonstrate that the proposed approach significantly improves the search performance and efficiency compared to MCTS. Multiple feasible routes have been successfully discovered and recognized by chemists, showcasing the practical utility of the method.
There are still some areas that need to be addressed in future research, for instance, in the encoding process, the proposed method converts continuous encoding search space into discrete search space for route searching. Although discrete encoding is effective in the current implementation using EA for retrosynthetic problems, there may be room for exploring more sophisticated encoding methods to further enhance search efficiency. Investigating alternative encoding techniques could potentially lead to improved performance and more efficient search processes.

### v2

@@ III Proposed method
This section initiates by framing the retrosynthetic problem as a tree search problem, where the objective is to find a sequence of routes leading from the root node to leaf nodes. Then the evolutionary algorithm efficiently deals with the tree search problem.
@@ III-A Retrosynthetic problem
where $R$ and $W$ represent reactants, $P$ represents product in Eq. 1a.
@@
Figure 4: K-cross search tree structure transformed by multi-step retrosynthesis process. Gray root node represents target product, $NN$ represents transformer model, and blue leaf node stands for raw material in the building block dataset.
@@ III-B Retrosynthetic optimization model
The number of potential chemical structures for organic molecules can reach an enormous scale, up to about $10^{60}$ compounds[48]. However, only a much smaller subset of compounds exhibits reasonable structural characteristics. Therefore, constructing a search space that explores the natural distribution of organic structures with structural rationality is a crucial challenge. The search space is defined as the top-k reactants with the highest inferred confidence in the single-step retrosynthetic reaction. The choice of the value $k$ is crucial as it impacts the search efficiency and the quality of the solutions. If the search space is too small, there is a risk of missing the global optimal solution. Conversely, if the search space is too large, the algorithm will consume more GPU memory and time, potentially leading to local optimality and reduced efficiency. Due to the vast search space, tree search alone may not be sufficient to effectively deal with the problem. Therefore, heuristic algorithms are employed to address this limitation and provide better solutions.
To tackle the retrosynthetic problem, it is converted into an optimization problem resembling tree search. Suppose that a multi-step retrosynthesis involves $q$ single-step retrosynthesis processes. $(x,r,\beta)$ represents a multi-step retrosynthesis, where $x=(x^{1},x^{2},\cdots,x^{q})$, and $x^{i}\in\{1,2,\cdots,k\}$ represents the $x^{i}$-th possible output of the $i$-th single-step retrosynthesis. Let $r=(r^{1},r^{2},\cdots,r^{q})$ denotes decoding results from $x$ using single-step retrosynthesis, as shown in Fig. 3, and $r^{i}$ represents SMILES expression (e.g., Fc1cccc(F)c1CBr.[N-]=[N+]=[N-]) of the $i$-th single-step retrosynthesis. $\beta=(\beta^{1},\beta^{2},\cdots,\beta^{q})$ denotes the corresponding probability of $r$. Take $(<2,2,1>,<r^{1},r^{2},r^{3}>,<0.8,0.9,0.7>)$ as an example, there are 3 steps: in the 1st step, $r^{1}$ is chosen as the second reactant of $NN$ with a probability of $\beta^{1}$=0.8; in the 2nd step, $r^{2}$ is chosen as the second reactant of $NN$ with a probability of $\beta^{2}$=0.9; in the 3rd step, $r^{3}$ is chosen as the first reactant of $NN$ with a probability of $\beta^{3}$=0.7 and it is a known molecule in $\Psi$.
@@ III-C2 Variation
The retrosynthetic problem is tackled using the Estimation of Distribution Algorithm (EDA) [32] during the execution of the EA process. EDA operator stands out from other operators as it generates new solutions by sampling from a histogram probabilistic model. This characteristic enables EDA to produce better or more similar solutions and converge faster. The basic framework of the proposed method is depicted in Fig. 6. The process of variation comprises the establishment of the probability model and sample. To construct the probability model, the variation space $[a^{i,0},a^{i,M}]$ for the $i$-th variable is divided into M bins, where $a^{i,0}=0$ and $a^{i,M}=1$ are the boundaries of the bins. Subsequently, $a^{i,1}$ and $a^{i,M-1}$ can be set as follows:
@@ IV Experiment
This section starts by introducing the datasets and parameter settings. It then conducts a comparative study using various methods and performs extensive experiments on four case products. Finally, several charts are provided along with corresponding analyses.
@@ IV-A Datasets
This experiment involved retrosynthetic route planning for four case products. Two commonly used benchmark datasets, namely, USPTO_50K [51, 52] and USPTO_MIT, were utilized to train single-step retrosynthetic model. The USPTO_50K dataset was extracted from the patent literature of the United States, while USPTO_MIT was previously employed by Lin et al.[3]. The distribution of reaction classes in the two datasets is presented in Table I. To define the terminal nodes or reactants, the building block dataset [3] is built, which contains 93563 commercially available molecules.
TABLE I: Descriptions of 10 reaction classes and the ratio of USPTO_50K and USPTO_MIT
@@
IV-B Parameters setting
During the experiment, the parameters are outlined below:
  • Population size. Set $N$ = 42. Larger populations typically have higher diversity, which can increase the robustness of the algorithm. But larger populations require more computational resources and time to evaluate and evolve each generation.
  • Thread pool size. Set $P$ = 42, which is equal to population size.
  • Termination condition. The population converges in about 150 generations at most; therefore, the search terminated after 200 iterations.
  • Number of bins. Set $M$ = 10. $M$ represents the number of bins in each dimension of the population, which ranges from 0 to 1. Because the results of the single-step model are encoded as the numbers 0-9, a total of 10 results. Smaller $M$ will miss some encoding results, while larger $M$ leads to computational redundancy.
Each experiment is executed for 30 independent runs, and the mean and variance are used for result analysis. The running program is written using python scripts. During the search process, the retrosynthetic routes are explored using 3 NVIDIA GeForce RTX 4090 GPUs.
TABLE II: Analysis of the number (mean(std.dev.))[percentage] of calling single-step model for MCTS and EA. Percentage in “[$\cdot$]” expresses increase or decrease of EA compared with MCTS in the same column and under the same product over 30 independent runs.
@@ IV-C Comparative experiment
The experimental section conducted a series of comparative experiments to validate the effectiveness of the proposed algorithm. In particular, it scrutinized the impact on the number of calls to the single-step model, the search capability in a larger search space, and the count of feasible search solutions. Additionally, several objective optimization functions were designed, and the algorithm’s search time was tested.
@@ IV-C1 Comparison experiment on calling single-step model
The frequency of calling single-step model directly affects both search time and search efficiency. MCTS and EA were employed to search for one, two, and three solutions for the four case products, as depicted in Table II. The values in the table indicate the average and standard deviation of the number of calling single-step model. As the number of solutions increase, both MCTS and EA require a higher count of model calls. However, EA consistently outperforms MCTS for one, two, and three solutions. Furthermore, across the four case products, the number of calling single-step model reduced by an average of 78.6%, 59.0%, and 53.9%, respectively. Consequently, EA exhibits faster search speed and higher search efficiency when compared to MCTS.
@@ IV-C2 Comparison experiment on the search capability
Tests with three different beam sizes (top-10, top-15, and top-20) were conducted to compare the search capabilities of the two algorithms in different search spaces.
Firstly, with the expansion of search space, the number of calling single-step model for both MCTS and EA increases significantly. As shown in Table IV(a), due to the advantage of EA search mechanism, under the beam size of top-10, top-15, and top-20, the number of calling single-step model of EA decreased by 53.9%, 27.6%, and 16.8% on average, respectively, compared to that of MCTS based on searching three feasible solutions.
Secondly, this experiment is conducted with a beam size of top-10. Notably, there is a progression in the complexity of molecular structures from DemoA to DemoD, which subsequently leads to longer search time. When focusing on the same target product, the search time for EA decreased significantly compared to that of MCTS, with an average reduction of 83.9% across the four case products based on searching three feasible solutions, as shown in Table IV(b).
Thirdly, as the search space increased in Table IV(c), the complexity of the search also rises, meanwhile, the number of search routes for four case products gradually decreased under the same number of iterations. However, EA outperforms MCTS in terms of the number of search routes across top-10, top-15, and top-20. Among them, the performance on top-15 and top-20 is normal, but the performance on top-10 is particularly noteworthy, with the number of search routes increasing 1.38 times on average. This could be attributed to the relatively simplistic and easily searchable molecular structure of DemoA.
@@
TABLE III: Analysis of search capability (mean(std.dev.))[percentage] on beam size top-10, top-15, top-20. Percentage in “[$\cdot$]” expresses increase or decrease of EA compared with MCTS in the same column and under the same product. Data in the table is over 30 independent runs.
@@
(a) Num of calling single-step model based on searching three feasible solutions.
@@
(b) Search time (sec) based on searching three feasible solutions.
@@
DemoA EA 9.53e+0(3.15e+0)[138$\uparrow$] 3.86e+0(1.01e+0)[18.0$\uparrow$] 2.71e+0(1.67e+0)[1.9$\uparrow$]
@@ IV-C3 Comparison experiment on different objective functions
$f_{*}$ $=-f(x_{i})$ (10)
or
$f_{\#}$ $=-e^{f(x_{i})}$ (11)
or
$f_{o}$ $=-\frac{f(x_{i})}{\sum_{j=1}^{n}f(x_{j})}$ (12)
During the selection process, solutions are ordered based on the fitness values for $f_{*}$, $f_{\#}$, and $f_{o}$. $f_{*}$ is the negation of the original objective function, and $f_{\#}$ engineered to increase the separation between target values. Meanwhile, $f_{o}$ utilizes the roulette wheel selection sampling method, inspired by the popular game of roulette[53]. This selection mechanism assigns a higher probability of selection to individuals with higher fitness values. Consequently, individuals with superior fitness have a greater chance of being chosen, leading to their retention in the evolutionary process. By favoring the selection of better individuals, this approach enhances the convergence and optimization capabilities of the genetic algorithm. Fig. 8 shows that $f_{*}$ yields more feasible solutions compared to the other objective functions, making it the most effective among them. To further assess the convergence of the three different objective functions on the four case products, the value range of $f_{\#}$ is normalized to match the same range as $f_{*}$ and $f_{o}$, which is from -1 to 0. As shown in Fig. 9, the results clearly demonstrate that $f_{*}$ outperforms the other two objective functions on DemoA, DemoB, and DemoC. Although $f_{o}$ shows some similarity to $f_{*}$ on DemoD, $f_{*}$ still performs better after 80-th iteration.
@@
Figure 9: Evolution of objective values of three different objective functions. Objective function $f_{*}$ outperforms the other two functions on the four target products.
@@ IV-D Visualization of results
In the retrosynthetic analysis, four case products were used. DemoA refers to Rufinamide[3], an antiepileptic triazole derivative. DemoB, DemoC, and DemoD are from Klucznik et al.[19], Li et al.[54], and Segler et al.[1], respectively. The retrosynthesis of these four products required 4, 2, 3, and 6 steps, respectively. As depicted in Fig. 11, DemoA’s search procedure comprises four steps leading to the final products being in $\Psi$. The impacted functional groups are highlighted in red. The digits before and after the “.” indicate the reaction type and the substance’s positioning in the results of the single-step model inference, correspondingly. These four case products have received approval from chemical experts.
During the search process, the distribution of the encoded population is visualized, as shown in Fig. 12. As the iteration increases, the population distributions gradually converge to the similar position. DemoA and DemoB demonstrate nearly rapid convergence after approximately 20 iterations, which can be attributed to their relatively simple molecular structures. In contrast, DemoC and DemoD, with more intricate molecular configurations, necessitate a broader range of reaction rules for analysis, presenting greater search challenges. As a result, these compounds converge around the 60-th iteration. Despite the intricate nature of the problem and the gradual rate of convergence, all four compounds ultimately achieve convergence.
@@
EA consistently outperforms MCTS in different aspects, as shown in Fig. 13. In the context of identifying three feasible solutions, there is a considerable decrease in the number of calling single-step model and search time. Additionally, what is particularly noteworthy is EA’s significant enhancement in searching feasible solutions under 200 iterations, especially for DemoA, because of its relatively simple molecular structure. Furthermore, the parallel implementation of EA results in an exponential reduction in time, making it a highly efficient approach for retrosynthetic route planning. Overall, the utilization of EA in retrosynthetic route planning demonstrates its capability to efficiently handle complex and challenging synthesis problems, thereby providing chemists with valuable and promising potential routes for target compounds.
Figure 13: The experiment of calling single-step model and search time is conducted based on identifying three feasible solutions. Additionally, the experiment of the number of feasible routes is carried out based on 200 iterations.
@@ V Conclusion
This work proposes a novel method for dealing with retrosynthetic route planning with Evolutionary Algorithms (EA). To be specific, a single-step model is used in the proposed method to learn the reaction rules from the datasets, and then a well-designed EA is adopted to deal with the optimization task, which is modeled from the retrosynthesis. The experiments show that the number of calling single-step model and the search time are reduced by and average of 53.9% and 83.9%, respectively. The number of search routes increases by 1.38 times. Multiple feasible routes have been successfully discovered and recognized by chemists, showcasing the practical utility of the method.
Although the method proposed in this study has demonstrated promising results, there are several directions that can be pursued:
  1. 1.
    Continuous encoding, used in the genetic operator, may affect the search efficiency of the algorithm. Therefore, selecting discrete encoding genetic operators can better match this research.
  2. 2.
    In practical applications, it may be more appropriate to model retrosynthetic route planning as a multi-objective optimization problem, especially in complex situations. By employing some promising generation paradigm [55] and solution estimation approach [56] that are specifically designed for multi-objective optimization, one can effectively address the challenges in such a problem.
  3. 3.
    After the implementation of the algorithm, it is necessary to continue to update some practical applications, such as a synthesis planning tool[57]. This could help chemists synthesize molecules more efficiently.
VI Acknowledgment
In this work, we would like to acknowledge the financial support from the National Natural Science Foundation of China (Grant 62306174), China Postdoctoral Science Foundation (Grant 2023M742259 and Grant 2023TQ0213), and Postdoctoral Fellowship Program of CPSF under Grant Number GZC20231588.